# Retrieval-Augmented Semantic Parsing: Improving Generalization with Lexical Knowledge

### v2

@@ Abstract
Open-domain semantic parsing remains a challenging task, as neural models often rely on heuristics and struggle to handle unseen concepts. In this paper, we investigate the potential of large language models (LLMs) for this task and introduce Retrieval-Augmented Semantic Parsing (RASP), a simple yet effective approach that integrates external symbolic knowledge into the parsing process. Our experiments not only show that LLMs outperform previous encoder-decoder baselines for semantic parsing, but that RASP further enhances their ability to predict unseen concepts, nearly doubling the performance of previous models on out-of-distribution concepts. These findings highlight the promise of leveraging large language models and retrieval mechanisms for robust and open-domain semantic parsing.
@@ 1 Introduction
Open-domain semantic parsing involves mapping natural language text to formal meaning representations, capturing the concepts, relations between them, and the contexts in which they appear Oepen and Lønning (2006); Hajič et al. (2012); Banarescu et al. (2013); Bos et al. (2017); Mart´ınez Lorenzo et al. (2022). Such meaning representations are applied in many downstream applications—ranging from database querying to embodied questions answering—where parsers must handle a vast array of concepts that may not appear in the training data. While neural encoder-decoder architectures have shown impressive performance in semantic parsing tasks, their reliance on training distributions constrains their ability to generalize, especially to out-of-distribution (OOD) concepts.
Most existing semantic parsers struggle to interpret the symbols, such as rare senses, often defaulting the unseen words to the most frequent meaning encountered during training. As a result, they fail to adapt to novel linguistic phenomena and remain limited to fixed patterns. Recent work Zhang et al. (2025) have attempted to mitigate these limitations by encoding concept representations numerically, forcing models to learn underlying symbolic knowledge from resources like WordNet Fellbaum (1998). However, these approaches require substantial preprocessing and intricate encodings that can be difficult for models to fully exploit.
In our work, instead, we explore the potential of large language models (LLMs), powerful decoder-only architectures with strong in-context learning capabilities and extensive pretraining, to enhance ability of semantic parsers to generalize. We pose two central research questions:
@@ 2.1 Discourse Representation Structure
Discourse Representation Theory (Kamp and Reyle, 1993, DRT) is a semantic modeling framework. The core component of DRT is the Discourse Representation Structure (DRS), a formal representation that captures the meaning of a discourse, which captures the essence of the text and covering linguistic phenomena like anaphors and temporal expressions. Unlike many other formalisms such as Abstract Meaning Representation (Banarescu et al., 2013, AMR) used for large-scale semantic annotation efforts, DRS covers logical negation, quantification, and discourse relations, has complete word sense disambiguation, and offers a language-neutral meaning representation. A Discourse Representation Structure (DRS) can be coded and visualised in various ways, which are all provided in Parallel Meaning Bank Abzianidze et al. (2017). In formal semantics they are often pictured in a human-readable box format. The clause notation was introduced to represent DRS in a sequential format suitable for machine learning models van Noord et al. (2018). To further simplify DRS, Bos (2023) proposed a variable-free format known as Sequence Box Notation (SBN). An example of the three different but logically equivalent formats is shown in Figure 1. Recent trends in using seq2seq models have led to a preference for sequence notation, which is also the format used in this paper.
Figure 1: Three formats of Discourse Representation Structure (DRS) for ”Mary didn’t commit a crime.”: the box notation, a directed acyclic graph, and the sequence notation.
@@ 2.3 Word Sense Disambiguation
In semantic parsing, WSD can be seen as a subtask Zhang et al. (2025), but it is more challenging because the parsing model must generate the correct sense directly without access to an explicitly provided sense inventory. However, traditional knowledge-based WSD offers a potential solution that inspires our approach: by retrieving and presenting all possible concepts as alternatives, we can explicitly provide external information to the model, thereby enhancing its generalization capability. As a consequence, this requires the model to be able to process long contexts, making the LLMs become the preferences, in particular retrieval augmented generation.
@@ 3 Retrieval-Augmented Semantic Parsing
Gold DRS female.n.02 Name ”Mary” time.n.08 TPR now birdwatch.v.01 Agent -2 Time -1 ELABORATION ¡1 female.n.02 ANA -3 see.v.01 Experiencer -1 Time +1 Stimulus +3 time.n.08 TPR now harrier.n.03 golden_eagle.n.01 entity.n.01 Sub -2 Sub -1 Sub +1 hobby.n.03
@@
$p_{decoder-only}(o|x)$ $=\prod^{n}_{i=1}p_{\theta}(o_{i}|x,o_{1:i-1})$ (1)
@@
$P_{{RASP}}(o|x$ $)=p(o^{\prime}|x)p_{decoder-only}(o|x,o^{\prime})$ (2)
$=p(o^{\prime}|x)\prod_{i=1}^{N}p_{\theta}(o_{i}|x,o^{\prime},o_{1:i-1})$
@@ 4.1 Datasets
To further assess the models’ ability to handle out-of-distribution (OOD) concepts, we adopt the challenge set proposed by Zhang et al. (2025), which is also derived from the PMB. Neural semantic parsers often default to the first sense of unknown concepts–an approach that can lead to ”lucky guesses” without truly understanding new words. The challenge set, consisting of 500 sentences, is deliberately designed to eliminate this shortcut. Each sentence includes at least one concept that does not appear in the training data and does not correspond to the first sense in the ontology. In total, the challenge set contains 410 unknown nouns, 128 verbs, and 65 modifiers (adjectives and adverbs). By evaluating on this set, we measure the true generalization capability of the models, testing whether they can correctly interpret novel concepts rather than relying on heuristic assigning.
@@ 4.2 Experiment Settings
Due to computational constraints, we select open-sourced LLMs with model sizes under 10B parameters, including phi3-4B, Mistral-7B, LLaMa3-3B, LLaMa3-8B, Gemma2-2B, Gemma2-9B, Qwen2.5-3B, and Qwen2.5-7B. These models strike a balance between state-of-the-art language understanding and manageable resource requirements. For fine-tuning, we employ Low-Rank Adaptation (Hu et al., 2021, LoRA), a parameter-efficient technique that introduces trainable low-rank matrices into the model’s layers, greatly reducing computational overhead.
@@ 5.1 Semantic Parsing on Standard Test
These higher scores are also reflected in Soft-SMatch and node-level F-scores, indicating that LLM-based models not only produce more structurally accurate meaning representations but also more reliably identify concept nodes. Additionally, Ill-Formed Rate (IFR) reductions suggest that these models generate fewer ill-structured outputs. In summary, these improvements highlight that large language models outperforms previous encoder-decoder models.
@@ 5.2 Performance on the Challenge Set
Table 4 provides the results on the challenge set, designed specifically to test the models’ ability to predict out-of-distribution (OOD) concepts. Here, we report Wu-Palmer similarities for unknown nouns, verbs, and modifiers (adjectives & adverbs). We calculate the Wu-Palmer similarities between the target concepts (out-of-distribution concepts) and the generated concepts (see examples in Table 5).
@@ 5.3 Error Analysis on the Challenge Set
The detective planted a bug in the suspect’s office to to gather evidence. plant.v.05 v.02 (0.22) v.05 (1.00)
@@
The visitors can observe various species of ray in the quarium. observe.v.02 v.01 (0.25) v.01 (0.25)
@@
We picked 22 instances, comprising 11 completely perfect predictions (WuP=1.00) and 11 imperfect predictions (WuP¡1.00) made by RASP, as presented in Table 5. With respect to the perfect predictions, it is evident that the retrieval significantly enhance the model’s ability to interpret most out-of-distribution concepts. For instance, in the text about birdwatching, the word ”hobby” clearly refers to a species of bird. The model without RAG defaults to the most frequent sense number, predicts hobby.n.01 (an auxiliary activity). In contrast, retrieval provides the glosses of each sense related to the noun ”hobby” and leads the model to pick hobby.n.03 (a falcon), by explicit lexical connections between ”falcon” in the gloss of hobby.n.03 and the context provided by ”birdwatching”.
@@ 6 Conclusion
This paper demonstrates that large language models (LLMs), even without retrieval augmentation, outperform previous encoder-decoder approaches in semantic parsing for Discourse Representation Structures, thereby answering our first research question in the affirmative, setting a new state of the art. We have also shown that our proposed Retrieval-Augmented Semantic Parsing (RASP) framework, which integrates external symbolic knowledge, further enhances the performance of LLMs. Notably, RASP nearly doubles the accuracy on out-of-distribution concepts, providing a direct answer to our second research question and confirming its robust generalization ability in open-domain scenarios. Our experiments show that by simply appending relevant information to the model input, the RASP approach offers a practical and intuitive approach that can be easily applied to other meaning representations used in natural language processing, such as AMR Banarescu et al. (2013) and BMR Mart´ınez Lorenzo et al. (2022).
@@ 7 Limitations
Another limitation is the retrieval source. Our implementation of RASP uses WordNet, so if a sense is not in WordNet, it will never be guessed. For example, ”velvet scooter” (a bird) is not in WordNet, nor is Cobb salad (a dish). Hence, RASP will never make a perfect prediction for such cases. Moreover, the glosses in WordNet, even though carefully crafted by lexicographers in most cases, are sometimes concise, lacking information to separate them from other senses. This makes it difficult for the models to accurately distinguish between different meanings (see Section 5.3).
@@ Appendix A Prompt
RASP Prompt user:
@@ Appendix B Experiment Settings
Phi-3-4B microsoft/Phi-3.5-mini-instruct

### v3

@@ Abstract
Open-domain semantic parsing remains a challenging task, as neural models often rely on heuristics and struggle to handle unseen concepts. In this paper, we investigate the potential of large language models (LLMs) for this task and introduce Retrieval-Augmented Semantic Parsing (RASP), a simple yet effective approach that integrates external lexical knowledge into the parsing process. Our experiments not only show that LLMs outperform previous encoder-decoder baselines for semantic parsing, but that RASP further enhances their ability to predict unseen concepts, nearly doubling the performance of previous models on out-of-distribution concepts. These findings highlight the promise of leveraging large language models and retrieval mechanisms for robust and open-domain semantic parsing.
@@ 1 Introduction
Open-domain semantic parsing involves mapping natural language text to formal meaning representations, capturing the concepts, relations between them, and the contexts in which they appear Oepen and Lønning (2006); Hajič et al. (2012); Banarescu et al. (2013); Bos et al. (2017); Mart´ınez Lorenzo et al. (2022). Such meaning representations are applied in many downstream applications—ranging from database querying to embodied question answering—where parsers must handle a vast array of concepts that may not appear in the training data. While neural encoder-decoder architectures have shown impressive performance in semantic parsing tasks, their reliance on training distributions constrains their ability to generalize, especially to out-of-distribution (OOD) concepts.
Most existing semantic parsers struggle to interpret the symbols, such as rare senses, often defaulting the unseen words to the most frequent meaning encountered during training. As a result, they fail to adapt to novel linguistic phenomena and remain limited to fixed patterns. Recent work Zhang et al. (2025) have attempted to mitigate these limitations by encoding concept representations symbolically, forcing models to learn underlying structural knowledge from resources like WordNet Fellbaum (1998). However, these approaches require substantial preprocessing and intricate encodings that can be difficult for models to fully exploit.
In our work, instead, we explore the potential of large language models, powerful decoder-only architectures with strong in-context learning capabilities and extensive pretraining, to enhance the ability of semantic parsers to generalize. We pose two central research questions:
@@ 2.1 Discourse Representation Structure
Discourse Representation Theory (Kamp and Reyle, 1993, DRT) is a semantic modeling framework. The core component of DRT is the Discourse Representation Structure (DRS), a formal representation that captures the meaning of a discourse, which captures the essence of the text and covers linguistic phenomena like anaphors and temporal expressions. Unlike many other formalisms such as Abstract Meaning Representation (Banarescu et al., 2013, AMR) used for large-scale semantic annotation efforts, DRS covers logical negation, quantification, and discourse relations. Moreover, DRS is equipped with complete word sense disambiguation, and offers a language-neutral meaning representation. A Discourse Representation Structure (DRS) can be coded and visualised in various ways, which are all provided in Parallel Meaning Bank Abzianidze et al. (2017). In formal semantics they are often pictured in a human-readable box format. The clause notation was introduced to represent DRS in a sequential format suitable for machine learning models van Noord et al. (2018). To further simplify DRS, Bos (2023) proposed a variable-free format known as Sequence Box Notation (SBN). An example of the three different but logically equivalent formats is shown in Figure 1. Recent trends in using seq2seq models have led to a preference for sequence notation, which is also the format used in this paper.
Figure 1: Three formats of Discourse Representation Structure (DRS) for ”Mary didn’t commit a crime.”: the box notation, a directed acyclic graph, and the sequence notation. These formalisms are mutually convertible without loss of information.
@@ 2.3 Word Sense Disambiguation
In semantic parsing, WSD can be seen as a subtask Zhang et al. (2025), but it is more challenging because the parsing model must generate the correct sense directly without access to an explicitly provided sense inventory. However, traditional knowledge-based WSD offers a potential solution that inspires our approach: by retrieving and presenting all possible concepts as alternatives, we can explicitly provide external information to the model, thereby enhancing its generalization capability. As a consequence, this requires the model to be able to process long contexts, making the LLMs be the preferred choice, in particular retrieval augmented generation.
@@ 3 Retrieval-Augmented Semantic Parsing
Gold DRS female.n.02 Name ”Mary” time.n.08 TPR now birdwatch.v.01 Agent -2 Time -1 ELABORATION <<1 female.n.02 ANA -3 see.v.01 Experiencer -1 Time +1 Stimulus +3 time.n.08 TPR now harrier.n.03 golden_eagle.n.01 entity.n.01 Sub -2 Sub -1 Sub +1 hobby.n.03
@@
$p_{\text{decoder-only}}(o\mid x)$ $=\prod_{i=1}^{n}p_{\theta}\!\left(o_{i}\mid x,\,o_{1:i-1}\right)$ (1)
@@
$p_{\text{RASP}}(o\mid x)$ $=p\!\left(o^{\prime}\mid x\right)\;p_{\text{decoder-only}}\!\left(o\mid x,\,o^{\prime}\right)$
$=p\!\left(o^{\prime}\mid x\right)\;\prod_{i=1}^{n}p_{\theta}\!\left(o_{i}\mid x,\,o^{\prime},\,o_{1:i-1}\right)$ (2)
@@ 4.1 Datasets
To further assess the models’ ability to handle out-of-distribution (OOD) concepts, we adopt the challenge set proposed by Zhang et al. (2025), which is also derived from the PMB. Neural semantic parsers often default to the first sense of unknown concepts–an approach that can lead to ”lucky guesses” without truly understanding new words. The challenge set, consisting of 500 sentences, is deliberately designed to eliminate this shortcut. Each sentence includes at least one concept that does not appear in the training data and does not correspond to the first sense in the ontology. In total, the challenge set contains 410 unknown nouns, 128 verbs, and 65 modifiers (adjectives and adverbs). By evaluating on this set, we measure the true generalization capability of the models, testing whether they can correctly interpret novel concepts rather than relying on heuristic assignment.
@@ 4.2 Experiment Settings
Due to computational constraints, we select open-sourced LLMs with model sizes under 10B parameters, including phi3-4B, Mistral-7B, LLaMa3.1-3B, LLaMa3.2-8B, Gemma2-2B, Gemma2-9B, Qwen2.5-3B, and Qwen2.5-7B. These models strike a balance between state-of-the-art language understanding and manageable resource requirements. For fine-tuning, we employ Low-Rank Adaptation (Hu et al., 2021, LoRA), a parameter-efficient technique that introduces trainable low-rank matrices into the model’s layers, greatly reducing computational overhead.
@@ 5.1 Semantic Parsing on Standard Test
These higher scores are also reflected in Soft-SMatch and node-level F-scores, indicating that LLM-based models not only produce more structurally accurate meaning representations but also more reliably identify concept nodes. Additionally, Ill-Formed Rate (IFR) reductions suggest that these models generate fewer ill-structured outputs. In summary, these improvements highlight that large language models outperform previous encoder-decoder models.
@@ 5.2 Performance on the Challenge Set
Table 4 provides the results on the challenge set, designed specifically to test the models’ ability to predict out-of-distribution (OOD) concepts. Here, we report Wu-Palmer similarities for unknown nouns, verbs, and modifiers (adjectives and adverbs). We calculate the Wu-Palmer similarities between the target concepts (out-of-distribution concepts) and the generated concepts (see examples in Table 5).
@@ 5.3 Error Analysis on the Challenge Set
The detective planted a bug in the suspect’s office to gather evidence. plant.v.05 v.02 (0.22) v.05 (1.00)
@@
The visitors can observe various species of ray in the aquarium. observe.v.02 v.01 (0.25) v.01 (0.25)
@@
We picked 22 instances, comprising 11 completely perfect predictions (WuP=1.00) and 11 imperfect predictions (WuP<<1.00) made by RASP, as presented in Table 5. With respect to the perfect predictions, it is evident that the retrieval significantly enhances the model’s ability to interpret most out-of-distribution concepts. For instance, in the text about birdwatching, the word ”hobby” clearly refers to a species of bird. The model without RAG defaults to the most frequent sense number, predicts hobby.n.01 (an auxiliary activity). In contrast, retrieval provides the glosses of each sense related to the noun ”hobby” and leads the model to pick hobby.n.03 (a falcon), by explicit lexical connections between ”falcon” in the gloss of hobby.n.03 and the context provided by ”birdwatching”.
@@ 6 Conclusion
This paper demonstrates that LLMs, even without retrieval augmentation, outperform previous encoder-decoder approaches in semantic parsing for Discourse Representation Structures, thereby answering our first research question in the affirmative, setting a new state of the art. We also show that our proposed Retrieval-Augmented Semantic Parsing (RASP) framework, which integrates external lexical knowledge, further enhances the performance of LLMs. Notably, RASP nearly doubles the accuracy on out-of-distribution concepts, which answers our second research question and confirms robust generalization ability of RASP in open-domain scenarios. Our experiments show that by simply appending relevant information to the model input, the RASP approach offers a practical and intuitive approach that can be easily applied to other meaning representations used in natural language processing, such as AMR Banarescu et al. (2013) and BMR Mart´ınez Lorenzo et al. (2022).
@@ 7 Limitations
Another limitation is the retrieval source. Our implementation of RASP uses WordNet, so if a sense is not in WordNet, it will never be guessed. For example, ”velvet scooter” (a bird) is not in WordNet, nor is Cobb salad (a dish). Hence, RASP will never make a perfect prediction for such cases. Moreover, the glosses in WordNet, even though carefully crafted by lexicographers in most cases, are sometimes concise, lacking information to separate them from other senses. This makes it difficult for the models to accurately distinguish between different meanings (see Section 5.3). For future work, the BabelNet, ConceptNet, or extended WordNet Delmonte et al. (2012); Navigli and Ponzetto (2012); Delmonte et al. (2015); Speer et al. (2017) can be considered as a better choice for concept in meaning representations.
@@ Appendix A Prompt
Normal Prompt user:
@@ Appendix B Experiment Settings
Phi3-4B microsoft/Phi-3.5-mini-instruct